test on electronic lamp manufactured by a company indicated their life span could be assumed to be normally distributed about a mean of 1750 hours with a standard derivation of 70 hours. calculate the probability that an electronic lamp purchased burn out: (1) in more than 2000 hours (2) between 1600 and 1900 hours.


1
Expert's answer
2021-10-04T02:27:45-0400

1)


"z^*=\frac{2000-1750}{70}=3.57\\P(Z>3.57)=0"

2)


"z^*_1=\frac{1600-1750}{70}=-2.14\\z^*_2=\frac{1900-1750}{70}=2.14\\P(-2.14<Z<2.14)=2(0.016)=0.032"